what is this in standard form slope is-2 and passes through the point(3,-5)
Answer by Alan3354(69443)

what is this in standard form slope is-2 and passes through the point(3,-5)
=========
Use y-y1 = m*(x-x1) where m=slope and (x1,y1) is the point.